OpenStudy (anonymous):

find the volume of ellipsoidal solid given by 4x^2+4y^2+z^2=16

OpenStudy (dumbcow):

rewrite in standard form divide equation by 16 \[\frac{x^2}{4} + \frac{y^2}{4} +\frac{z^2}{16} = 1\] http://en.wikipedia.org/wiki/Ellipsoid#Volume a = 2, b=2, c = 4

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\int\limits_{0}^{2}\int\limits_{0}^{4}4x^2+4y^2+z^2dzdydx\]

OpenStudy (anonymous):

is that true

OpenStudy (anonymous):

@dumbcow